# A New Model-Independent Precision Test of General Relativity using LISA Bright Standard Sirens

### v1

@@ 6.1 Case with the fixed Hubble constant
The correlation between the EM wave luminosity distance at a specific redshift (z) and critical cosmological parameters, such as the BAO scale ($\theta_{\text{BAO}}$) and the sound horizon ($r_{s}$), is succinctly expressed by the equation
$\mathcal{F}(z)=\frac{d_{l}^{\rm GW}(z)\theta_{\rm BAO}(z)}{(1+z)r_{s}}.$ (6.1)
@@
This equation is central to the analysis of the frictional term, linking it to three significant cosmological lengths: (i)GW Luminosity Distance: The distance for detectable GWs events is estimated through the parameter estimation of GWs events as described in Subsection 3.3. (ii)BAO Scale ($\theta_{\text{BAO}}$): Derived from the 2PACF and expressed in Equation 5.1, this scale is utilized in a versatile manner, allowing for application across various surveys. (iii)Sound Horizon Distance: Estimated by analyzing CMB acoustic oscillations through missions like WMAP [63], Planck [64], ACTPol [166], SPT-3G [167], and CMB-S4 [168]. The first peak in the angular power spectrum at recombination reflects the sound horizon scale, measured precisely by Planck as $147.09\pm 0.26$ Mpc at the drag epoch and redshift: Inferred from the EM counterparts.
@@
To estimate the posterior distribution of $\mathcal{F}(z)$ as a function of redshift for $n_{\text{GW}}$ GWs sources, we employ a Hierarchical Bayesian framework. The posterior distribution is represented as follows [56]
@@
$\times$ $d\theta_{BAO}^{i}P(\theta_{BAO}^{i})\mathcal{L}(d_{l}^{GW^{i}}|\mathcal{F}(z^{i}),\theta_{BAO}^{i},r_{s},z^{i}).$
In this framework, $P(\mathcal{F}(z)|\{d^{\text{GW}}\},\{z\})$ denotes the posterior probability density of $\mathcal{F}(z)$, given the GWs data $d^{\text{GW}}$ and the spectroscopic redshifts $\{z\}$ of the EM counterparts for $i$ events. The BAO angular scale measurements $\theta_{\text{BAO}}$ at each redshift of the GW sources, where an EM counterpart can be detected through a galaxy survey, are marginalized. The term $\mathcal{L}(d_{l}^{\text{GW}^{i}}|\mathcal{F}(z^{i}),\theta_{\text{BAO}}^{i},r_{s},z^{i})$ corresponds to the likelihood function, with $P(r_{s})$ and $P(\theta_{\text{BAO}}^{i})$ representing the prior probabilities for the sound horizon and BAO scale, respectively. $\Pi(\mathcal{F}(z))$ indicates the prior distribution for $\mathcal{F}(z)$, reflecting pre-existing knowledge or assumptions about the frictional term before the data is considered. A flat prior from 0.1 to 2 is adopted for $\mathcal{F}(z)$. This investigation focuses solely on bright sirens, assuming that the redshift of the host galaxies of the GW sources is measured spectroscopically. This implies a precise or assumed concentration of redshift values for these sources.
6.2 Case with varying the Hubble constant
In previous discussions, we concentrated on analyzing the frictional term, denoted as $\mathcal{F}(z)$ in Equation 6.1, under the assumption of a constant Hubble constant ($H_{0}$). Moving forward, we expand our analysis to concurrently estimate both $\mathcal{F}(z)$ and $H_{0}$. This approach allows for the joint estimation of the frictional term and the Hubble constant, leveraging constant sound horizon measurements across redshifts to normalize $\mathcal{F}(z)$. Our primary objective remains the precise determination of $\mathcal{F}(z)$, while concurrently assessing $H_{0}$. This dual estimation strategy is especially pertinent with the advent of future detectors, which will enhance our capability to probe deeper into redshifts, thus enabling either combined or independent measurements of these parameters. This methodology is particularly promising for addressing the so-called Hubble tension (for a comprehensive review [169]), the notable discrepancy between the locally measured values of $H_{0}$ and those inferred from CMB observations [170].
By employing a hierarchical Bayesian framework, we efficiently estimate these parameters using a comprehensive array of observational data, encompassing GWs detections, CMB data, and large-scale structure surveys, as elaborated in earlier sections. The hierarchical integration of these datasets is meticulously designed to manage their distinct uncertainties and biases effectively. This integration produces correlated constraints that significantly enhance our understanding of cosmological dynamics, help resolve discrepancies among observational datasets, and uncover critical relationships between key cosmological parameters.
The joint posterior distribution for $\mathcal{F}(z)$ and $H_{0}$, based on these considerations, is formulated as follows
$P(\mathcal{F}(z),H_{0}|\{d^{GW}\},\{z\})\propto\Pi(\mathcal{F}(z))\Pi(H_{0})\prod_{i=1}^{n_{GW}}\iiint$ (6.3)
$\times$ $dr_{s}P(r_{s}^{i})dd_{l}^{GW^{i}}d\theta_{BAO}^{i}P(\theta_{BAO}^{i})\mathcal{L}(d_{l}^{GW^{i}}|\mathcal{F}(z^{i}),H_{0},\theta_{BAO}^{i},r_{s},z^{i}).$
In this model, $P(\mathcal{F}(z),H_{0}|{d^{\text{GW}}},{z})$ denotes the joint posterior probability density function for $\mathcal{F}(z)$ and $H_{0}$, incorporating the GWs data $d^{\text{GW}}$ and spectroscopic redshifts ${z}$ for each event. The likelihood function is represented by $\mathcal{L}(d_{l}^{\text{GW}^{i}}|\mathcal{F}(z)^{i},H_{0}^{i},\theta_{\text{BAO}^{i}},r_{s},z^{i})$. The terms $\Pi(\mathcal{F}(z))$ and $\Pi(H_{0})$ define the prior distributions for $\mathcal{F}(z)$ and $H_{0}$, respectively, encapsulating pre-existing beliefs or assumptions prior to analyzing the data. Within our hierarchical Bayesian framework, we utilize flat priors for both parameters, indicating minimal initial bias. The range for $\mathcal{F}(z)$ is set from 0.1 to 2.0, and for $H_{0}$, from 40 km/s/Mpc to 100 km/s/Mpc.
7 Results
This study exclusively examines bright sirens detectable by the space-based GWs observatory LISA, focusing on events with a redshift up to $z=6$ across multiple population models. We present the frictional term, $\mathcal{F}(z)$, measurements under two scenarios: with a fixed Hubble constant ($H_{0}$) and with a varying $H_{0}$. The comparative results are illustrated in Figure 10 and Figure 11. Specifically, Figure 10 showcases the impact of single-event measurements, and Figure 11 highlights how event detection rates and population assumptions influence the precision of $\mathcal{F}(z)$ measurements.
Figure 10: This violin plot depicts the posterior distribution of the non-GR parameter $\mathcal{F}(z)$, illustrating its variation with cosmic redshift for LISA across both constant and varying $H_{0}$ scenarios for a single event. The upper x-axis shows the corresponding 1-$\sigma$ error bars for each measurement. Beyond redshift $z=3.0$, the plot uses a fainter color to indicate that the BAO scale cannot be measured beyond this redshift with currently planned galaxy surveys. A line at $\mathcal{F}(z)=1$ is included to represent the GR case. This figure underscores the enhanced capability of the future space-based detector LISA to measure $\mathcal{F}(z)$ with increased precision across extended cosmic distances.
Figure 11: This plot illustrates the precision in reconstructing the redshift evolution of the Planck mass, represented by $\mathcal{F}(z)$, for the Light Seed, Heavy Seed, and Heavy Seed+delay models using the upcoming space-based LISA observatory. The observation time is 4 years for the Light Seed and Heavy Seed models and 10 years for the Heavy Seed+delay model, assuming that EM counterparts are detected for 75% of the events. The plot highlights LISA’s accuracy in measuring the function $\mathcal{F}(z)$ up to a redshift of 6. Beyond a redshift of 3.0, a fainter color is used to indicate that the BAO scale cannot be measured beyond this redshift with the currently planned galaxy surveys.
7.1 Reconstruction of F(z)
In cases where $H_{0}$ is held constant, it acts as an overall normalization factor for $\mathcal{F}(z)$ across different redshifts. This setup provides a unique opportunity to normalize $\mathcal{F}(z)$ measurements to 1 if an independent and precise measurement of $\mathcal{F}(z)$ in our local universe is possible. Such normalization, particularly effective at low redshifts, helps eliminate the dependence on $H_{0}$ and demonstrates the potential deviations from GR due to the cumulative effects of modified gravity wave propagation over cosmological distances. The uncertainty in the measurement of the frictional term, $\mathcal{F}(z)$, is inversely proportional to the square root of the number of GW sources, denoted as $\mathrm{1/\sqrt{N_{GW}}}$. This relationship suggests that increasing the count of GW sources detected by LISA enhances the precision of $\mathcal{F}(z)$ measurements. However, the accuracy improvement reaches a plateau beyond a certain number of sources, primarily due to errors associated with the BAO scale. The precision in estimating the BAO scale emerges as a critical factor limiting the refinement of $\mathcal{F}(z)$ estimation accuracy. Thus, advancing the accuracy of BAO scale measurements is essential for further enhancements in determining the frictional term with redshift. Essentially, optimizing the number of GW observations, alongside improvements in BAO scale accuracy and GW luminosity distance measurements, is crucial for advancing our understanding of $\mathcal{F}(z)$ and probing new aspects of fundamental physics.
On the GWs front, the accuracy of $\mathcal{F}(z)$ relies predominantly on the precision of the luminosity distance to GWs sources, $D_{l}^{GW}(z)$. Enhancements in measuring $D_{l}^{GW}(z)$ would significantly boost the fidelity of $\mathcal{F}(z)$ estimations. Given the degeneracy of $D_{l}^{GW}(z)$ with other GWs parameters, notably the inclination angle ($i$), acquiring more accurate measurements of the inclination angle, potentially through EM counterparts, could substantially reduce uncertainties in $D_{l}^{GW}(z)$ and, by extension, refine $\mathcal{F}(z)$ estimates [171]. The effect of peculiar velocities on redshift inference from EM counterparts is notably critical at lower redshifts ($z<0.03$) but becomes negligible at higher redshifts [172]. Given that the majority of GW sources detected by LISA are expected to be at higher redshifts, this effect is disregarded in our analysis.
7.2 Reconstruction of F(z) and Hubble Constant
Building upon the hierarchical Bayesian framework introduced earlier to estimate $\mathcal{F}(z)$ with a fixed $H_{0}$, we now extend this approach to concurrently measure $\mathcal{F}(z)$ and the Hubble constant, $H_{0}$. In this extended framework, we derive the EM luminosity distance by integrating various cosmological scales, including the BAO scale, sound horizon distance, and redshift. Figures 10 and 11 provide a detailed analysis of $\mathcal{F}(z)$ measurements under scenarios of both fixed and varying $H_{0}$. Significantly, when $H_{0}$ is allowed to vary, the uncertainties in measuring $\mathcal{F}(z)$ increase by approximately a factor of 2 to 2.5, relative to scenarios with a fixed $H_{0}$. In Figure 11, we specifically illustrate the improved precision in tracking the redshift-dependent variations of the Planck mass ($\mathcal{F}(z)$) for MBBHs systems, emphasizing enhanced measurements in the redshift range from 1.0 to 6.0. Additionally, Figure 12 illustrates the optimal combined measurements of $\mathcal{F}(z)$ and the Hubble constant ($H_{0}$).
In Figure 13, we show the measurement of $H_{0}$ marginalized over $\mathcal{F}(z)$ for the three different population models, for all events up to a redshift of $z=6.0$. The accuracy of $\mathcal{F}(z)$ estimates relies heavily on precise measurements of the BAO scale and $D_{l}^{GW}(z)$. Improved precision in BAO scale readings directly bolsters the accuracy of $\mathcal{F}(z)$ estimates. On the GW front, the fidelity of $\mathcal{F}(z)$ assessments is primarily dependent on the accuracy of luminosity distances, $D_{l}^{GW}(z)$. Enhancements in measuring $D_{l}^{GW}(z)$ substantially improve $\mathcal{F}(z)$ evaluations, as previously discussed in the previous subsection.
Figure 12: This plot illustrates the optimal combined measurements of $F(z)$ and the Hubble constant ($H_{0}$) for the Light Seed, Heavy Seed, and Heavy Seed+delay models using the forthcoming space-based LISA observatory. The observation periods are 4 years for the Light Seed and Heavy Seed models, and 10 years for the Heavy Seed+delay model, assuming that EM counterparts are detected for 75% of the events. These optimal measurements occur at redshifts 3.2, 4.0, and 2.4 for the Light Seed, Heavy Seed, and Heavy Seed+delay models, respectively.
Figure 13: This plot shows the posterior distribution of the Hubble constant ($H_{0}$) marginalized over $F(z)$ for all events up to a redshift of 6, based on three models: Light Seed, Heavy Seed, and Heavy Seed+delay, using the upcoming space-based LISA observatory. The observation periods are 4 years for both the Light Seed and Heavy Seed models and 10 years for the Heavy Seed+delay model, assuming that EM counterparts are detected for 75% of the events. The uncertainties in the measurements are provided in parentheses in the legend.

### v2

@@ 6.1 Case with the fixed Hubble constant
The correlation between the EM luminosity distance at a specific redshift (z) and critical cosmological parameters, such as the BAO scale ($\mathrm{\theta_{BAO}}$) and the sound horizon ($\mathrm{r_{s}}$), is succinctly expressed by the Equation 2.7. This equation is central to the analysis of the frictional term, linking it to three significant cosmological lengths: (i)GW Luminosity Distance: The distance for detectable GW events is estimated through the parameter estimation of GW events as described in Subsection 3.3. (ii)BAO Scale ($\mathrm{\theta_{BAO}}$): Derived from the 2PACF and expressed in Equation 5.1, this scale is utilized in a versatile manner, allowing for application across various surveys. (iii)Sound Horizon Distance: Estimated by analyzing CMB acoustic oscillations through missions like WMAP [63], Planck [64], ACTPol [169], SPT-3G [170], and CMB-S4 [171]. The first peak in the angular power spectrum at recombination reflects the sound horizon scale, measured precisely by Planck as $\mathrm{147.09\pm 0.26}$ Mpc at the drag epoch and redshift: Inferred from the EM counterparts.
@@
To estimate the posterior distribution of $\mathrm{\mathcal{F}(z)}$ as a function of redshift for $\mathrm{n_{GW}}$ GW sources, we employ a Hierarchical Bayesian framework. The posterior distribution is represented as follows [57]
$\mathrm{P(\mathcal{F}(z))\propto\Pi(\mathcal{F}(z))\prod_{i=1}^{n_{GW}}\iiint dr_{s}P(r_{s})dD_{L}^{GW^{i}}d\theta_{BAO}^{i}P(\theta_{BAO}^{i})\mathcal{L}(D_{L}^{GW^{i}}|\mathcal{F}(z^{i}),\theta_{BAO}^{i},r_{s},z^{i})}.$ (6.1)
In this framework, $\mathrm{P(\mathcal{F}(z))}$ denotes the posterior probability density of $\mathrm{\mathcal{F}(z)}$. The term $\mathrm{\mathcal{L}(D_{L}^{\text{GW}^{i}}|\mathcal{F}(z^{i}),\theta_{\text{BAO}}^{i},r_{s},z^{i})}$ corresponds to the likelihood function, which represents the posterior distribution of the GW luminosity distance. In this expression, $\mathrm{P(r_{s})}$ and $\mathrm{P(\theta_{BAO}^{i})}$ representing the prior probabilities for the sound horizon and BAO scale, respectively. $\mathrm{\Pi(\mathcal{F}(z))}$ indicates the prior distribution for $\mathrm{\mathcal{F}(z)}$, reflecting pre-existing knowledge or assumptions about the frictional term before the data is considered. A flat prior from 0.1 to 2 is adopted for $\mathrm{\mathcal{F}(z)}$. This investigation focuses solely on bright sirens, assuming that the redshift of the host galaxies of the GW sources is measured spectroscopically. This assumption implies precise measurements of redshift for these sources, as the error in redshift measurements is very small compared to the errors associated with the BAO scale and distance measurements.
6.2 Case with varying the Hubble constant
In previous discussions, we concentrated on analyzing the frictional term, denoted as $\mathrm{\mathcal{F}(z)}$ in Equation 2.7, under the assumption of a constant Hubble constant ($\mathrm{H_{0}}$). Moving forward, we expand our analysis to concurrently estimate both $\mathrm{\mathcal{F}(z)}$ and $\mathrm{H_{0}}$. Our primary objective remains the precise determination of $\mathrm{\mathcal{F}(z)}$, while concurrently assessing $\mathrm{H_{0}}$. This dual estimation strategy is especially pertinent with the advent of future detectors, which will enhance our capability to probe deeper into redshifts, thus enabling either combined or independent measurements of these parameters. This methodology is particularly promising for addressing the so-called Hubble tension (for a comprehensive review [172]), the notable discrepancy between the locally measured values of $\mathrm{H_{0}}$ and those inferred from CMB observations [173].
By employing a hierarchical Bayesian framework, we efficiently estimate these parameters using a comprehensive array of observational data, encompassing GW detections, CMB data, and large-scale structure surveys, as elaborated in earlier sections. The hierarchical integration of these datasets is meticulously designed to manage their distinct uncertainties and biases effectively. This integration produces correlated constraints that significantly enhance our understanding of cosmological dynamics, help resolve discrepancies among observational datasets, and uncover critical relationships between key cosmological parameters. The joint posterior distribution for $\mathrm{\mathcal{F}(z)}$ and $\mathrm{H_{0}}$, based on these considerations, is formulated as follows
$\mathrm{P(\mathcal{F}(z),H_{0})\propto\Pi(\mathcal{F}(z))\Pi(H_{0})\prod_{i=1}^{n_{GW}}\iiint}\mathrm{dr_{s}P(r_{s}^{i})dD_{L}^{GW^{i}}d\theta_{BAO}^{i}P(\theta_{BAO}^{i})\mathcal{L}(D_{L}^{GW^{i}}|\mathcal{F}(z^{i}),H_{0},\theta_{BAO}^{i},r_{s},z^{i})}.$ (6.2)
In this model, $\mathrm{P(\mathcal{F}(z),H_{0}})$ denotes the joint posterior probability density function for $\mathrm{\mathcal{F}(z)}$ and $\mathrm{H_{0}}$. The likelihood function is represented by $\mathrm{\mathcal{L}(D_{L}^{\text{GW}^{i}}|\mathcal{F}(z)^{i},H_{0}^{i},\theta_{\text{BAO}^{i}},r_{s},z^{i})}$. The terms $\mathrm{\Pi(\mathcal{F}(z))}$ and $\mathrm{\Pi(H_{0})}$ define the prior distributions for $\mathrm{\mathcal{F}(z)}$ and $\mathrm{H_{0}}$, respectively, encapsulating pre-existing beliefs or assumptions prior to analyzing the data. Within our hierarchical Bayesian framework, we utilize flat priors for both parameters, indicating minimal initial bias. The range for $\mathrm{\mathcal{F}(z)}$ is set from 0.1 to 2.0, and for $\mathrm{H_{0}}$, from 40 km/s/Mpc to 100 km/s/Mpc.
7 Results
This study exclusively examines bright sirens detectable by the space-based GWs observatory LISA, focusing on events with a redshift up to $\mathrm{z=6}$ across multiple population models. We present the frictional term, $\mathrm{\mathcal{F}(z)}$, measurements under two scenarios: with a fixed Hubble constant ($\mathrm{H_{0}}$) and with a varying $\mathrm{H_{0}}$. The comparative results are illustrated in Figure 11 and Figure 12. Specifically, Figure 11 showcases the impact of single-event measurements, and Figure 12 highlights how event detection rates and population assumptions influence the precision of $\mathrm{\mathcal{F}(z)}$ measurements.
Figure 11: This violin plot depicts the posterior distribution of the non-GR parameter $\mathrm{\mathcal{F}(z)}$, illustrating its variation with cosmic redshift for LISA across both constant and varying $\mathrm{H_{0}}$ scenarios for a single event. The upper x-axis shows the corresponding 1-$\mathrm{\sigma}$ error bars for each measurement. Beyond redshift $\mathrm{z=3.0}$, the plot uses a fainter color to indicate that the BAO scale cannot be measured beyond this redshift with currently planned galaxy surveys. A line at $\mathrm{\mathcal{F}(z)=1}$ is included to represent the GR case. This figure underscores the enhanced capability of the future space-based detector LISA to measure $\mathrm{\mathcal{F}(z)}$ with increased precision across extended cosmic distances. The plot is based on a single MBBH merger event with a fixed total mass of $\mathrm{10^{6}}$ $\mathrm{M_{\odot}}$ having a mass ratio $q=1$.
Figure 12: This plot illustrates the precision in reconstructing the redshift evolution of the Planck mass, represented by $\mathcal{F}(z)$, for the Light Seed, Heavy Seed, and Heavy Seed+delay models using the upcoming space-based LISA observatory. The observation time is 4 years for the Light Seed and Heavy Seed models and 10 years for the Heavy Seed+delay model. The filled regions represent the $1\sigma$ intervals for the assumption of $75\%$ EM counterpart detection, while the error bars with lines indicate the $1\sigma$ intervals for the assumption of $50\%$ EM counterpart detection case. The plot highlights LISA’s accuracy in measuring the function $\mathcal{F}(z)$ up to a redshift of 6. Beyond a redshift of 3.0, a fainter color is used to indicate that the BAO scale cannot be measured beyond this redshift with the currently planned galaxy surveys.
7.1 Reconstruction of F(z)
In cases where $\mathrm{H_{0}}$ is held constant, it acts as an overall normalization factor for $\mathrm{\mathcal{F}(z)}$ across different redshifts. This setup provides a unique opportunity to normalize $\mathrm{\mathcal{F}(z)}$ measurements to 1 if an independent and precise measurement of $\mathcal{F}(z)$ in our local universe is possible. Such normalization, particularly effective at low redshifts, helps eliminate the dependence on $\mathrm{H_{0}}$ and demonstrates the potential deviations from GR due to the cumulative effects of modified gravity wave propagation over cosmological distances. The uncertainty in the measurement of the frictional term, $\mathrm{\mathcal{F}(z)}$, is inversely proportional to the square root of the number of GW sources, denoted as $\mathrm{1/\sqrt{N_{GW}}}$. This relationship suggests that increasing the count of GW sources detected by LISA enhances the precision of $\mathrm{\mathcal{F}(z)}$ measurements. However, the accuracy improvement reaches a plateau beyond a certain number of sources, primarily due to errors associated with the BAO scale. The precision in estimating the BAO scale emerges as a critical factor limiting the refinement of $\mathrm{\mathcal{F}(z)}$ estimation accuracy. Thus, advancing the accuracy of BAO scale measurements is essential for further enhancements in determining the frictional term with redshift. Essentially, optimizing the number of GW observations, alongside improvements in BAO scale accuracy and GWs luminosity distance measurements, is crucial for advancing our understanding of $\mathrm{\mathcal{F}(z)}$ and probing new aspects of fundamental physics.
On the GWs front, the accuracy of $\mathrm{\mathcal{F}(z)}$ relies predominantly on the precision of the luminosity distance to GWs sources, $\mathrm{D_{L}^{GW}(z)}$. Enhancements in measuring $\mathrm{D_{L}^{GW}(z)}$ would significantly boost the fidelity of $\mathrm{\mathcal{F}(z)}$ estimations. Given the degeneracy of $\mathrm{D_{L}^{GW}(z)}$ with other GWs parameters, notably the inclination angle ($\mathrm{i}$), acquiring more accurate measurements of the inclination angle, potentially through EM counterparts, could substantially reduce uncertainties in $\mathrm{D_{L}^{GW}(z)}$ and, by extension, refine $\mathrm{\mathcal{F}(z)}$ estimates [174]. The effect of peculiar velocities on redshift inference from EM counterparts is notably critical at lower redshifts ($\mathrm{z<0.03}$) but becomes negligible at higher redshifts [175]. Given that the majority of GW sources detected by LISA are expected to be at higher redshifts, this effect is disregarded in our analysis.
7.2 Reconstruction of F(z) and Hubble Constant
Building upon the hierarchical Bayesian framework introduced earlier to estimate $\mathrm{\mathcal{F}(z)}$ with a fixed $\mathrm{H_{0}}$, we now extend this approach to concurrently measure $\mathrm{\mathcal{F}(z)}$ and the Hubble constant, $\mathrm{H_{0}}$. In this extended framework, we derive the EM luminosity distance by integrating various cosmological scales, including the BAO scale, sound horizon distance, and redshift. Figures 11 provide a detailed analysis of $\mathrm{\mathcal{F}(z)}$ measurements under scenarios of both fixed and varying $\mathrm{H_{0}}$ with only six events up to $z=6$. The uncertainty on the parameter $\mathrm{\mathcal{F}(z)}$ increased by approximately a factor of 2 to 2.5 when $\mathrm{H_{0}}$ is allowed to vary relative to scenarios with a fixed $\mathrm{H_{0}}$. This scenario with only six events corresponds to a pessimistic case with only $4\%$ EM bright events for the Light Seed model and the Heavy Seed model and about $30\%$ for the Heavy Seed+Delay model in four years of observation period. This corresponds to a single events at every redshift up to $z=6$. We have shown the constraints for a few scenarios of detectable bright sirens from LISA in appendix A. If the efficiency of LISA bright sirens are less than this, then one needs to consider a longer observation period of the LISA mission for at least one detection of bright standard siren.
In Figure 12, we specifically illustrate the improved precision in tracking the redshift-dependent variations of the Planck mass ($\mathrm{\mathcal{F}(z)}$) for MBBHs systems, emphasizing enhanced measurements in the redshift range from $z=1.2$ to $z=6.0$. Additionally, Figure 13 illustrates the optimal joint measurements of $\mathrm{\mathcal{F}(z)}$ and the Hubble constant ($\mathrm{H_{0}}$). In Figure 14, we show the measurement of $H_{0}$ marginalized over $\mathrm{\mathcal{F}(z)}$ for the three different population models, for all events up to a redshift of $\mathrm{z=6.0}$. The results on $\mathrm{\mathcal{F}(z)}$ and $H_{0}$ for scenario with MBBHs distribution having a mass ratio beyond $q=1$ is shown in the appendix B.
Figure 13: This plot illustrates the optimal combined measurements of $\mathrm{\mathcal{F}(z)}$ and the Hubble constant ($\mathrm{H_{0}}$) for the Light Seed, Heavy Seed, and Heavy Seed+delay models using the forthcoming space-based LISA observatory. The observation periods are 4 years for the Light Seed and Heavy Seed models, and 10 years for the Heavy Seed+delay model, assuming that EM counterparts are detected for $\mathrm{75\%}$ of the events. These optimal measurements occur at redshifts 3.2, 4.0, and 2.4 for the Light Seed, Heavy Seed, and Heavy Seed+delay models, respectively.